OpenStudy (pphalke):

nico is saving money for his college education. he invests some money at 6% and $1900 less than the amount at 5%. the investments produced a total of $191 interest in 1 yr. how much did he invest at each rate? i have a deadline i need to complete this, please someone help me. i need it bad

OpenStudy (photon336):

didn't someone answer this?

OpenStudy (pphalke):

they did but the wrong percents were there its 6% and 5%

OpenStudy (photon336):

hold on let me see

OpenStudy (pphalke):

ok

OpenStudy (photon336):

@pphalke do you know how to solve this?

OpenStudy (pphalke):

no i dont and it has to be for 6% and 5% 2 different rates

OpenStudy (pphalke):

he invests some money at 6% and $1900 less than the amount at 5%. the investments produced a total of $191 interest in 1 yr. how much did he invest at each rate? at each rate is what the answer needs to be

OpenStudy (pphalke):

$9600 =x

OpenStudy (pphalke):

now what do i do @Photon336

OpenStudy (photon336):

Hold on I need to correct something

OpenStudy (pphalke):

ok

OpenStudy (photon336):

the investments produced a total of $191 interest in 1 yr. how much did he invest at each rate? he invests some money at 6% translates to 0.06x $1900 less than the amount at 5%. so less than means subtraction so it becomes (1900-x) and we take 5% of the total, so we multiply the whole thing by 0.05 0.05(x-1900) \[0.06x+0.05(x-1900) = 191\] now we get 0.06+0.05 = This should be the correct expression can you solve this now? 0.11x-1900 =191 0.11x = 2091

OpenStudy (pphalke):

@Photon336 i get a weird number when i divide by 0.11x it comes out 19009.09